# Rua Dr. Bento Teobaldo Ferraz 271, 01140-070, São Paulo, SP, Brazil

### v3

@@ 1 Introduction
    When the symmetry is ungauged, we classify the currents in all non-trivial representations of $S_{N}$ and show that those below spin 10 acquire anomalous dimensions. We further argue that this should remain true at large but finite $m$.
@@ 2.2 Anomalous dimensions of singlet currents
The initial observation is that when the operator content of the theory changes continuously, taking $g_{\sigma}\to 0$ in (16) leaves behind a short multiplet with $(h,\bar{h})=(J,0)$ and a long multiplet with $(h,\bar{h})=(J,1)$. For a given spin $J$, it is conceptually straightforward to enumerate a basis of $(h,\bar{h})=(J,0)$ operators whose elements will be denoted $T^{K}$. These are what we have been referring to as (UV) currents. It is similarly possible to produce a basis for the space of $(h,\bar{h})=(J,1)$ operators with elements $V^{K}$. In practice, one should regard these bases as arbitrary which means they are almost certainly not nice enough to yield $T^{K}$ and $V^{K}$ operators which match up one-to-one. As such, the divergence of a given $T^{K}$ at one loop is a non-trivial linear combination of the $V^{K}$ whose coefficients are encoded in the all-important matrix $b^{K}_{\;\;L}$. Before this linear combination is taken, the $V^{K}$ will be referred to as candidates for being the divergence of a spin-$J$ current. We will also employ the phrase potential divergences, dropping the word “potential” for brevity when the context makes this clear. [^12] Since we are studying the lifting induced by $\sigma$, one more condition we will often impose is that four copies of $\phi_{(1,2)}$ must be present in the expression for $V^{K}$ to make it a divergence candidate or a potential divergence.
@@
Setting (18) equal to (19), we can solve for $b^{K}_{\;\;L}$ by performing a calculation at one-loop instead of two. Moreover, the three-point function at zero loops that we are instructed to integrate vanishes at separated points due to ¯$\bar{\partial}T^{K}(z_{1})$. This leads to the appearance of a delta function which makes the integral trivial.
Due to the simplification of this integral, one can observe a key fact which is that the only way for $T^{K}$ to stay conserved is to have $b^{K}_{\;\;L}=0$ for all $L$. This can only happen if the three-point functions in (18) similarly vanish for all $L$. While $b^{K}_{\;\;L}\neq 0$ for a single $L$ is sufficient to conclude that $T^{K}$ is not a dilation eigenstate with weights of $(h,\bar{h})=(J,0)$ in the IR, we are more interested in establishing this for arbitrary linear combinations of the $T^{K}$. As such, the check to undertake is that $b^{K}_{\;\;L}$ is a matrix whose rank is equal to the number of rows. This is what led to the conclusion in Antunes:2022vtb that $S_{N}$ singlets up to spin-10 for $N=5,6,7$ all lift. The next two sections are about applying the same strategy to non-singlets.
@@ 10 $(N(N(36+3N)+81)-74)N(N-1)/120$ 89 226 481
Table 2: List of degeneracies for all diagonal-Virasoro primary currents in the tensor product theory.
@@ 3.3 Refined partition functions
$Z_{(4)}(\tau,\bar{\tau})$ $=\frac{1}{4!}\left[Z(\tau,\bar{\tau})^{4}+6Z(2\tau,2\bar{\tau})Z(\tau,\bar{\tau})^{2}+3Z(2\tau,2\bar{\tau})^{2}+8Z(3\tau,3\bar{\tau})Z(q,\bar{q})+6Z(4\tau,4\bar{\tau})\right]\,,$
@@ 3.4 Comments on finite $m$
A reasonable worry about the analysis of Antunes:2022vtb is whether the lifting of currents is expected to persist when the spins $J$ of the currents start scaling with some positive power of $m$. Indeed, conformal perturbation theory presumably breaks down when this happens. We claim that the physics at sufficiently large but finite $m$ should not be modified substantially when compared to the large-$m$ expansion. Indeed, the fundamental mechanism that ensures the lifting of currents is the overwhelming growth of potential divergences with which they can recombine. As we showed in the subsection above, at fixed $J$ the degeneracies of divergences $d^{\prime}_{J,1}$ when compared to the degeneracies of currents $d_{J,0}$ scale as
@@
When we take $m$ to be finite, the spectrum is modified by the appearance of additional null states. This reduces the number of potential divergences but also of initial currents. A simple exercise in Virasoro characters and counting shows that this effect has negligible consequences on the ability for the currents to recombine with their divergence candidates. We recall that the for the $m$’th minimal model, the Virasoro character for a doubly-degenerate primary reads
@@ 4 Broken conservation and the fate of the infrared
We will now explain how to turn the above results into a concrete algorithm for strengthening the evidence that coupled minimal model fixed points are irrational. This leads to the main result of this paper which is that currents of spin less than $10$ outside the $\widehat{\mathfrak{Vir}}$ identity multiplet all lift for $N=5,6,7$. The steps below will be similar to those that were used (and explained very tersely) in Antunes:2022vtb but without any assumptions about the currents or divergences being singlets of $S_{N}$.
@@ 4.1 Construction of operators
Table 8: The degeneracies of spin $J\leq 10$ currents which are relevant for the first step of our algorithm for constructing operators. They have been refined by $S_{N}$ representation under the assumption that $N\gg 1$. These can be predicted by following the finite $N$ techniques of section 3 by raising $N$ until the numbers stabilize.
@@
where the sum runs over tuples $(i_{1},\dots,i_{a})$ of distinct elements. In the notation of Antunes:2022vtb , this is written as
@@ 4.2 Lifting from three-point functions
$z_{12}^{2J-1}z_{13}z_{23}\bar{z}_{23}^{2}\left<T^{K}(z_{1})V^{L}(z_{2},\bar{z}_{2})\sigma(z_{3},\bar{z}_{3})\right>$ (65)
@@
$L_{-n_{1}}\dots L_{-n_{b}}\phi(z_{2},\bar{z}_{2})\phi(z_{3},\bar{z}_{3})\right>&=\mathcal{L}_{-n_{1}}\left<L_{-n_{2}}\dots L_{-n_{b}}\phi(z_{2},\bar{z}_{2})\phi(z_{3},\bar{z}_{3})\right>\\ \mathcal{L}_{n_{1}}&\equiv\frac{n_{1}-1}{4(z_{3}-z_{2})^{n_{1}}}-\frac{\partial_{3}}{(z_{3}-z_{2})^{n_{1}-1}}\end{split}$ (67)
@@
This concludes the list of operations that must be carried out in this approach to verify that UV currents lift. The arXiv submission for the latest version of this paper includes Python code for this purpose. It is sufficient to verify that anomalous dimensions are non-zero for the cases we have tested which are $J<10$ and $N=5,6,7$.
@@ 6 Outlook
and do brute force searches for fixed points of the couplings $a_{ijkl}$ and $b_{i}$ in the spirit of os20 or attempt to constrain their general properties in the spirit of rs18 . Additional parameters can be introduced by taking slightly different large $m$ limits for each copy, i.e we can take $m_{i}=m\gamma_{i}$ for different values of the constants $\gamma_{i}$. Furthermore, it is also possible to consider other classes of exactly solvable rational conformal field theories to be coupled in the UV. A particularly convenient set are minimal models of ADE type $\mathcal{W}$-algebras. In this case a classification of weakly relevant couplings is possible, but additional OPE coefficients must be determined to write down the beta-function equations. In this context, additional Virasoro primary currents are present because of the $\mathcal{W}$-symmetry. A relevant question is whether a diagonal copy of this symmetry can be preserved or whether only the Virasoro algebra survives. For example, it would be interesting to understand what the fate of diagonal $\mathcal{W}_{3}$ symmetry is in the coupled Potts models studied in djlp98 .
@@ Acknowledgements.
We are grateful for conversations with Christopher Beem, Nathan Benjamin, Jasper Jacobsen, Elias Kiritsis, Dalimil Mazáč, Hirosi Ooguri, Eric Perlmutter, Sylvain Ribault, Junchen Rong, Slava Rychkov, Adar Sharon, Yifan Wang and Jingxiang Wu. AA received funding from the German Research Foundation DFG under Germany’s Excellence Strategy – EXC 2121 Quantum Universe – 390833306 and is funded by the European Union (ERC, FUNBOOTS, project number 101043588). Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council Executive Agency. Neither the European Union nor the granting authority can be held responsible for them. CB received funding from the São Paulo Research Foundation (FAPESP) grants 2019/24277-8, 2021/14335-0 and 2023/03825-2. AA thanks IPhT Saclay where part of this work was presented. CB thanks the Perimeter Institute for hospitality during the final stages of this work. Research at the Perimeter Institute is supported in part by the Government of Canada through NSERC and by the Province of Ontario through MRI.

### v5

@@ 1 Introduction
    When the symmetry is ungauged, making use of refined partition functions, we classify the currents in all non-trivial representations of $S_{N}$ and show that those below spin 10 acquire anomalous dimensions. We further argue that this should remain true at large but finite $m$.
@@ 2.2 Anomalous dimensions of singlet currents
The initial observation is that when the operator content of the theory changes continuously, taking $g_{\sigma}\to 0$ in (16) leaves behind a short multiplet with $(h,\bar{h})=(J,0)$ and a long multiplet with $(h,\bar{h})=(J,1)$. For a given spin $J$, it is conceptually straightforward to enumerate a basis of $(h,\bar{h})=(J,0)$ operators whose elements will be denoted $T^{K}$. These are what we have been referring to as (UV) currents. It is similarly possible to produce a basis for the space of $(h,\bar{h})=(J,1)$ operators with elements $V^{K}$. In practice, one should regard these bases as arbitrary which means they are almost certainly not nice enough to yield $T^{K}$ and $V^{K}$ operators which match up one-to-one. As such, the divergence of a given $T^{K}$ at one loop is a non-trivial linear combination of the $V^{K}$ whose coefficients are encoded in the all-important matrix $b^{K}_{\;\;L}$. Before this linear combination is taken, the $V^{K}$ will be referred to as divergence candidates of a spin-$J$ current. [^12] Since we are studying the lifting induced by $\sigma$, one more condition we will often impose is that four copies of $\phi_{(1,2)}$ must be present in the expression for $V^{K}$ to make it a divergence candidate.
@@
Setting (18) equal to (19), we can solve for $b^{K}_{\;\;L}$ by performing a calculation at one-loop instead of two. Moreover, the three-point function at zero loops that we are instructed to integrate vanishes at separated points due to ¯$\bar{\partial}T^{K}(z_{1})$. This leads to the appearance of a delta function which makes the integral trivial, yielding the relation
$b^{K}_{L}=\pi C^{K}_{L\,\sigma}+O(1/m)\,,$ (20)
where $C^{K}_{L\,\sigma}$ is the OPE coefficient for $T^{K}(z_{1})V^{L}(z_{2},\bar{z}_{2})\sigma(z_{3},\bar{z}_{3})\right>$. Due to the simplification of this quantity, one can observe a key fact which is that the only way for $T^{K}$ to stay conserved is to have $b^{K}_{\;\;L}=C^{K}_{L\,\sigma}=0$ for all $L$. While $b^{K}_{\;\;L}\neq 0$ for a single $L$ is sufficient to conclude that $T^{K}$ is not a dilation eigenstate with weights of $(h,\bar{h})=(J,0)$ in the IR, we are more interested in establishing this for arbitrary linear combinations of the $T^{K}$. As such, the check to undertake is that $b^{K}_{\;\;L}$ is a matrix whose rank is equal to the number of rows. This is what led to the conclusion in Antunes:2022vtb that $S_{N}$ singlets up to spin-10 for $N=5,6,7$ all lift. The next two sections are about applying the same strategy to non-singlets.
@@ 10 $(N(N(36+3N)+81)-74)N(N-1)/120$ 89 226 481
Table 2: List of degeneracies for all diagonal-Virasoro primary currents in the tensor product theory. Note that $J=10$ has been included here even though lifting for some irreps has only been checked up to $J=9$.
@@ 3.3 Refined partition functions
$Z_{(4)}(\tau,\bar{\tau})$ $=\frac{1}{4!}\left[Z(\tau,\bar{\tau})^{4}+6Z(2\tau,2\bar{\tau})Z(\tau,\bar{\tau})^{2}+3Z(2\tau,2\bar{\tau})^{2}+8Z(3\tau,3\bar{\tau})Z(\tau,\bar{\tau})+6Z(4\tau,4\bar{\tau})\right]\,,$
@@ 3.4 Comments on finite $m$
A reasonable worry about the analysis of Antunes:2022vtb is whether the lifting of currents is expected to persist when the spins $J$ of the currents start scaling with some positive power of $m$. Indeed, conformal perturbation theory presumably breaks down when this happens. We claim that the physics at sufficiently large but finite $m$ should not be modified substantially when compared to the large-$m$ expansion. Indeed, the fundamental mechanism that ensures the lifting of currents is the overwhelming growth of divergence candidate with which they can recombine. As we showed in the subsection above, at fixed $J$ the degeneracies of divergences $d^{\prime}_{J,1}$ when compared to the degeneracies of currents $d_{J,0}$ scale as
@@
When we take $m$ to be finite, the spectrum is modified by the appearance of additional null states. This reduces the number of divergence candidate but also of initial currents. A simple exercise in Virasoro characters and counting shows that this effect has negligible consequences on the ability for the currents to recombine with their divergence candidates. We recall that the for the $m$’th minimal model, the Virasoro character for a doubly-degenerate primary reads
@@ 4 Broken conservation and the fate of the infrared
[…]
@@ 4.1 Construction of operators
Table 8: The degeneracies of spin $J\leq 10$ currents which are relevant for the first step of our algorithm for constructing operators. They have been refined by $S_{N}$ representation under the assumption that $N\gg 1$. These can be predicted by following the finite $N$ techniques of section 3 by raising $N$ until the numbers stabilize. Note that the final analysis has been limited to $J\leq 9$ due to memory constraints.
@@
where the sum runs over tuples $(i_{1},\dots,i_{a})$ of distinct elements. [^19] As a typographical note, the subscript of the Virasoro charge is $n_{j,k}$. We need both $j$ and $k$ because there are multiple copies and potentially multiple Virasoro charges being multiplied for each copy. In the notation of Antunes:2022vtb , this is written as
@@ 4.2 Lifting from three-point functions
$C^{K}_{L\,\sigma}=z_{12}^{2J-1}z_{13}z_{23}\bar{z}_{23}^{2}\left<T^{K}(z_{1})V^{L}(z_{2},\bar{z}_{2})\sigma(z_{3},\bar{z}_{3})\right>$ (66)
@@
$L_{-n_{1}}\dots L_{-n_{b}}\phi(z_{2},\bar{z}_{2})\phi(z_{3},\bar{z}_{3})\right>&=\mathcal{L}_{-n_{1}}\left<L_{-n_{2}}\dots L_{-n_{b}}\phi(z_{2},\bar{z}_{2})\phi(z_{3},\bar{z}_{3})\right>\\ \mathcal{L}_{-n_{1}}&\equiv\frac{n_{1}-1}{4(z_{3}-z_{2})^{n_{1}}}-\frac{\partial_{3}}{(z_{3}-z_{2})^{n_{1}-1}}\end{split}$ (68)
@@
This concludes the list of operations that must be carried out in this approach to verify that UV currents lift. The arXiv submission for the latest version of this paper includes Python code for this purpose. It is sufficient to verify that anomalous dimensions are non-zero for the cases we have tested which are $J<10$ and $N=5,6,7$. It is worth emphasizing that these results are exact since the entries of (66) have been computed using rational numbers. With more resources, one could avoid having to stop at $N=7$ (or indeed any fixed value) by working with rational functions of $N$.
@@ 6 Outlook
and do brute force searches for fixed points of the couplings $a_{ijkl}$ and $b_{i}$ in the spirit of os20 or attempt to constrain their general properties in the spirit of rs18 . Additional parameters can be introduced by taking slightly different large $m$ limits for each copy, i.e we can take $m_{i}=m\gamma_{i}$ for different values of the constants $\gamma_{i}$. Furthermore, it is also possible to consider other classes of exactly solvable rational conformal field theories to be coupled in the UV. A straightforward example is to take the D-series minimal models, which also admit the $\sigma$ and $\epsilon$ deformations. Working at large but finite $m$, our analysis goes through since the same set of currents and divergence candidates is present. However, analytic continuation between the D-series minimal models appears to be more challenging. One therefore needs to keep in mind potential (but unlikely) cancellations between terms of different orders in the large-$m$ expansion. A richer generalization can be obtained by coupling minimal models of ADE type $\mathcal{W}$-algebras. In this case a classification of weakly relevant couplings is possible, but additional OPE coefficients must be determined to write down the beta-function equations. In this context, additional Virasoro primary currents are present because of the $\mathcal{W}$-symmetry. A relevant question is whether a diagonal copy of this symmetry can be preserved or whether only the Virasoro algebra survives. For example, it would be interesting to understand what the fate of diagonal $\mathcal{W}_{3}$ symmetry is in the coupled Potts models studied in djlp98 .
@@ Acknowledgements.
We are grateful for conversations with Christopher Beem, Nathan Benjamin, Jasper Jacobsen, Elias Kiritsis, Dalimil Mazáč, Hirosi Ooguri, Eric Perlmutter, Jiaxin Qiao, Sylvain Ribault, Junchen Rong, Slava Rychkov, Adar Sharon, Yifan Wang and Jingxiang Wu. AA received funding from the German Research Foundation DFG under Germany’s Excellence Strategy – EXC 2121 Quantum Universe – 390833306 and is funded by the European Union (ERC, FUNBOOTS, project number 101043588). Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council Executive Agency. Neither the European Union nor the granting authority can be held responsible for them. CB received funding from the São Paulo Research Foundation (FAPESP) grants 2019/24277-8, 2021/14335-0 and 2023/03825-2. AA thanks IPhT Saclay where part of this work was presented. CB thanks the Perimeter Institute for hospitality during the final stages of this work. Research at the Perimeter Institute is supported in part by the Government of Canada through NSERC and by the Province of Ontario through MRI.